OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. E(-2, 2), F(5, 1)

OpenStudy (anonymous):

first use slope point form to determine the equation of a line that is , y-y1 = m (x-x1) where m = slope = (y2-y1)/(x2-x1) (x1,y1) and (x2,y2) = points through which the line passes

OpenStudy (anonymous):

so how is the equation written (answers please)

OpenStudy (anonymous):

m = (1-2)/{5-(-)2} = -1/7 therefore y-1 = -1/7*(x-5)

OpenStudy (anonymous):

you sure?

OpenStudy (anonymous):

it cant have fractions

OpenStudy (anonymous):

or decimals

OpenStudy (anonymous):

now u have to convert above into a standard form multiply 7 in both sides

OpenStudy (anonymous):

\[y-1 = -\frac{ (x-5) }{ 7 }\] multiply 7 both sides

OpenStudy (anonymous):

then rearrange the equation in such a manner that u get this form Ax +By = C (standard form)

OpenStudy (anonymous):

@kavin_pitre is it clear to you?

OpenStudy (anonymous):

could you do the work i dont have a calculator and yes it is clear but will you solve it for me please

OpenStudy (anonymous):

i know now how to solve a problem like this now thank you

OpenStudy (anonymous):

:) cheers

OpenStudy (anonymous):

lol